What is the net force on a 1,000 kg object accelerating at 3 m/s2?
1
Expert's answer
2020-11-23T10:26:10-0500

By the definition of the Newton's Second Law of Motion, we have:


"F=ma=1000\ kg\cdot 3\ \dfrac{m}{s^2}=3000\ N."

Answer:

"F=3000\ N."